Hi all, I'm having trouble finding the determinant of the following matrix A.
Let Matrix A =
(2 -1 3)
(1 4 4)
(1 0 2)
Any help would be most appreciated.
Thanks, -Nick.

Answer by mathmontrealmath(2)

use the diagonal method, repeat the first two columns to the right of the matrix:
(2 -1 3)2 -1
(1 4 4)1 4
(1 0 2)1 0
then multiply numbers diagonally starting from the top left:
2.4.2-1.4.1+3.1.0
then - (multiply numbers diagonally starting from the bottom left)
1.4.3-0.4.2-2.1.(-1)
add these answers, you get the determinant value:
2.4.2-1.4.1+3.1.0 = 16-4+0=12
1.4.3-0.4.2-2.1.(-1)=12-0+2=14
add: 12+14=26
done